Determine the work done by a force 

F--> = xy î+ yz ĵ+ xzk̂

in taking a particle along 

the path defined by the equation

r(t) = tî + 2t²ĵ + t³k̂e from t=0 to t=1 . Is force conservative?


1
Expert's answer
2021-01-30T06:25:31-0500

Given force is "\vec{F} = xy î+ yz ĵ+ xzk̂"

Since "\vec{r(}t) = tî + 2t²ĵ + t³k̂"


x is changing according to the equation "x = t"

y is changing according to the equation "y=2t^2"

z is changing according to the equation "z = t^3"

Putting value of x,y and z

Now, F will be, "\vec{F} = 2t^3 î+ 2t^5 ĵ+ t^4k̂"

"d\vec{r} = dt\hat{i}+4tdt\hat{j}+3t^2dt\hat{k}"


Then work done by the object is given by,

"W = \int_0^1 \vec{F} .d\vec{r} = \int_0^1(2t^3 î+ 2t^5 ĵ+ t^4k̂).(dt\hat{i}+4tdt\hat{j}+3t^2dt\hat{k})"


"W = \int_0^1 (2t^3+11t^6)dt = [ \frac{2}{4}t^4+\frac{11}{7}t^7 ]_0^1 = \frac{1}{2}+\frac{11}{7} = \frac{29}{14}J"